### v1

# Systematic Constraint Formulation and Collision Free Trajectory Planning Using Space-Time Graphs of Convex Sets

Abstract
In this paper, we create optimal, collision-free, time-dependent trajectories through cluttered dynamic environments. The many spatial and temporal constraints make finding an initial guess for a numerical solver difficult. Graphs of Convex Sets (GCS) and the recently developed Space-Time Graphs of Convex Sets formulation (ST-GCS) enable us to generate optimal minimum distance collision-free trajectories without providing an initial guess to the solver. We also explore the derivation of general GCS-compatible constraints and document an intuitive strategy for adapting general constraints to the framework. We show that ST-GCS produces results that are equivalent to the standard GCS formulation when the environment is static. We then show ST-GCS operating in dynamic environments to find minimum distance collision-free trajectories.
I Introduction
Figure 1: A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$ (a). Safe regions of space and time are partitioned into a Graph of Convex Sets (GCS), and then a collision-free trajectory is generated through the graph (b).
A fundamental problem in autonomous systems is generating a safe reference path or trajectory for a feedback controller to track. Numerous planning methods exist for computing geometric paths that satisfy task-specific objectives [1]. However, such paths are typically insufficient in dynamic or time-varying environments [2, 3], which require explicit modeling of how obstacles or goals change over time. In these cases, a trajectory—defined as a path parameterized over time—is more appropriate, as it enables temporal coordination and optimization in response to environmental changes [4, 5].
In motion planning, environments are frequently represented as discrete graphs by partitioning space into regions that enable efficient search-based methods. Algorithms such as Dijkstra’s Algorithm [6], A*, and D*[7] are widely used to compute globally optimal paths without requiring an initial guess[8]. However, the computational complexity of graph search algorithms scales with graph size [9, 10], and the resulting paths often neglect physical constraints such as continuity or smoothness. As a result, discrete solutions typically require post-processing to convert them into feasible continuous trajectories [11], which may cause deviation from global optimality.
Continuous optimization methods, by contrast, can directly enforce physical constraints on trajectories. These methods model obstacles using analytic functions such as signed distance fields [12], control barrier functions [13], or polygonal representations [14], and they optimize a cost function subject to these constraints. However, a key limitation of continuous optimization is its reliance on a good initial guess. In complex or cluttered environments, poor initializations can lead to suboptimal or infeasible solutions. To mitigate this, sampling-based methods are often used to generate feasible initial trajectories [15, 16]. Still, these approaches are especially sensitive to moving obstacles, which introduce additional temporal nonlinearities and local minima [17, 18].
To overcome these limitations, Marcucci et al. introduced the Graphs of Convex Sets (GCS) framework [19, 20, 21]. GCS is a trajectory optimization approach that integrates the discrete structure of graph search with the constraint handling capabilities of continuous optimization. The framework constructs a graph of convex regions and formulates trajectory optimization as a combined graph search and convex optimization program. Notably, GCS has been applied to shortest-path planning [20, 22], minimum-time navigation in maze-like environments [21], and robotic arm motion planning in high-dimensional configuration spaces [23, 24, 25].
A key strength of GCS is that it finds continuous trajectories without requiring an initial guess. It uses a branch-and-bound algorithm to iteratively refine candidate solutions until a globally optimal trajectory is found. Additionally, GCS provides metrics that indicate the quality of candidate solutions, allowing rapid selection of high-quality local optima [22].
Despite these advantages, the current GCS pipeline has notable limitations. As explained in [21], it separates spatial and temporal planning by first computing a globally optimal collision-free path through a static environment, then determining the trajectory timing through an external nonlinear optimization. This latter step not only sacrifices global optimality guarantees but can also fail to produce a valid solution even when one exists. Moreover, existing approaches provide limited treatment of dynamic obstacle avoidance, which must be further developed and integrated before this framework can be reliably applied to safe trajectory planning.
A particularly challenging aspect contributing to this complexity is the uniform formulation of constraints across the entire graph as formulated in  [19]. Even basic constraints, such as enforcing continuity between spline segments, must be expressed in a highly specific mathematical form, which is often abbreviated or omitted entirely. This requirement increases implementation challenges compared to other frameworks. To our knowledge, prior work has not provided explicit, accessible formulations of these constraints, further hindering adoption. This paper seeks to address this gap by presenting a detailed methodology for formulating and implementing such constraints, thereby facilitating broader use of the GCS framework.
One notable attempt at obstacle avoidance is by Werner et al.[26], who recompute the graph of convex sets and regenerate the trajectory at each timestep. To accelerate this process, their method requires GPU-accelerated graph computation. However, it does not incorporate any prior knowledge of obstacle motion into the planning, and still relies on a nonlinear optimization step to apply timing to the path—necessitating a good initial guess. In contrast, Marcucci et al.[21] propose joint planning for two robotic arms by concatenating their configuration spaces. While effective for a small number of agents, this approach scales poorly due to exponential growth in the combined configuration space, limiting its applicability. Additionally, because agents often cannot control the motion of other objects in real-world scenarios, such methods are unsuitable for general-purpose dynamic obstacle avoidance.
The most promising obstacle avoidance GCS formulation to date is that of Tang et al. [27], who recently introduced the Space–Time Graph of Convex Sets (ST-GCS) framework for multi-robot motion planning. In this formulation, the configuration space is augmented with a time dimension, enabling constant-velocity obstacles to be represented in a manner compatible with the GCS framework. Their work considers only the other agents involved in joint planning between initial and final points in space. The method prioritizes minimum-time solutions and incorporates a mechanism to “reserve” space–time trajectories as dynamic obstacles for subsequent agents.
A limitation of this approach is that the current GCS framework requires that only one convex set in the graph may intersect the final trajectory state. As a result, Tang et al. conservatively select the final set, potentially preventing the framework from finding true minimum time trajectories even if they exist. Additionally, their formulation is restricted to planning piecewise-constant velocity trajectories and is not applied to higher-order splines. Critical constraints that ensure causality within the space–time configuration space are present, but their importance is not discussed or elaborated.
In this paper, we generate collision-free, minimum-distance trajectories with a fixed end time using the ST-GCS framework, enabling direct comparison with the standard GCS formulation in static settings. The minimum distance cost function allows us to validate the framework’s applicability to dynamic scenarios and assess its performance and limitations within time-varying environments. Figure 1 illustrates the ST-GCS formulation, depicting a moving two-dimensional object and a graph of convex sets generated in the time-augmented configuration space.
This paper makes two main contributions. First, we provide a systematic methodology, accompanied by step-by-step examples, for rewriting common constraints in the form required by the GCS framework. Second, building on the ST-GCS framework, we investigate the generation of minimum-distance, collision-free trajectories in dynamic 2D environments. By assuming known obstacle positions and constant obstacle velocities, our approach produces higher-order optimal spline trajectories without relying on nonlinear solvers, initial temporal guesses, or incurring unnecessary computational complexity.
The remainder of this paper is structured as follows. Section II provides an overview of the underlying mathematics for graphs of convex sets and Bézier Splines. Section III details our methodology and presents examples of formulating GCS-compatible constraints. Section IV presents the results of optimizing trajectories with GCS in dynamic environments. Finally, Section V concludes the paper and discusses directions for future work.
II Background
The GCS optimization framework provides strong guarantees for the end user, and it is rooted in advanced concepts from convex analysis and optimization theory. In this section, we will highlight the key concepts for understanding GCS at a high level. For a more comprehensive understanding of GCS, we refer readers to [19, 20, 21, 22, 23, 24, 25, 26, 28].
II-A Convex Tools
As the name suggests, convex sets are fundamental to the GCS framework. In this subsection, we introduce key definitions and equations regarding convex sets, but refer readers to chapters two and three of [28] for a more comprehensive treatment of the topic.
A set $\mathcal{S}~\subseteq~\mathbb{R}^{n}$ is convex if, for any two points within the set, the straight line segment connecting them lies entirely within $\mathcal{S}$. Convexity can be demonstrated by taking two points $a,b\in\mathcal{S}$ and showing that the linear interpolation between the two always remains in the set $\mathcal{S}$. Formally, this is expressed as:
$(1-\lambda)a+\lambda b\in\mathcal{S}\quad\forall a,b\in\mathcal{S}\quad\lambda\in[0,1].$ (1)
Often, sets we encounter in practice are not convex. The convex hull of set $\mathcal{S}$, denoted as conv($\mathcal{S}$), is the smallest convex set that contains $\mathcal{S}$. An illustration of this is shown in Figure 2. If $\mathcal{S}$ is already convex, then conv($\mathcal{S}$) = $\mathcal{S}$.
Another important concept in convex optimization is that of a cone. A set $\mathcal{S}$ is called a cone (or said to be non-negative homogeneous) if for all $x\in\mathcal{S}$ and $\lambda\geq 0$ we have $\lambda x\in\mathcal{S}$ (shown in Figure 3). A set is a convex cone if it is both convex and a cone.
Notably, a set can be a cone without being convex. For example, consider the union of the first and third quadrants of $\mathbb{R}^{2}$. This set is a cone because any vector in the first and third quadrants of $\mathbb{R}^{2}$ can be scaled by a nonnegative number and remain within the set, but it is not convex. We will exclusively use convex cones in this paper.
Figure 2: A nonconvex set $\mathcal{S}$ shown in (a) and its associated convex hull $\text{conv}(\mathcal{S})$ shown in (b).
Figure 3: If $\mathcal{S}$ is a cone, the vector $x\in\mathcal{S}$ will still be in $\mathcal{S}$ after scaling by a non-negative constant.
Convex combinations are another useful tool we will use when translating constraints into the GCS framework. A convex combination $x_{c}$ of points $\vec{x}_{1},\vec{x}_{2},...\vec{x}_{m}\in\mathbb{R}^{n}$ is defined as
$\vec{x}_{c}=\sum_{i=1}^{m}\vec{x}_{i}\lambda_{i}\quad\text{where}\quad\lambda_{i}\geq 0,\quad\sum_{i=1}^{m}\lambda_{i}=1.$ (2)
This guarantees that $\vec{x}_{c}$ lies within $\text{conv}\left(\{\vec{x}_{1},\vec{x}_{2},...,\vec{x}_{m}\}\right)$. A linear combination of nonnegative coefficients can be normalized into a convex combination by dividing by the sum of the coefficients, assuming that the sum is positive. Specifically, if
$\lambda\vec{x}_{c}=\sum_{i=1}^{m}\vec{x}_{i}\lambda_{i}\quad\text{where}\quad\lambda_{i}\geq 0,\quad\lambda=\sum_{i=1}^{m}\lambda_{i}$ (3)
then
$\vec{x}_{c}$ $=\frac{1}{\lambda}\sum_{i=1}^{m}\vec{x}_{i}\lambda_{i}\quad\text{where}$ (4)
$\lambda_{i}$ $\geq 0,\quad\lambda>0,\quad\frac{1}{\lambda}\sum_{i=1}^{m}\lambda_{i}=1.$
This relationship will be helpful when we derive and interpret the GCS-compatible continuity and differentiability constraints in the methods section.
Another important tool used in the GCS framework is the closure of the homogenization of a closed set. Given a closed set $\mathcal{S}\in\mathbb{R}^{n}$, the closure of the homogenization of the set, denoted $\text{clh}(\mathcal{S})$, allows us to construct a cone in $\mathbb{R}^{n+1}$ (see Figure 4). The $\text{clh}()$ operation is defined as
$\text{ clh}(\mathcal{S})$= $\left\{\text{concat}(\vec{x},\lambda)\in\mathbb{R}^{n+1}:\lambda\geq 0,\vec{x}\in\lambda\mathcal{S}\right\},$ (5)
where $\text{concat}(\vec{x},\vec{v})$ concatenates vectors $\vec{x}\in\mathbb{R}^{n}$ and $\vec{v}\in\mathbb{R}^{m}$ into $\mathbb{R}^{n+m}$. The original set can be recovered by taking a cross-section at $\lambda=1$. If $\mathcal{S}$ is convex, then $\text{clh}(\mathcal{S})$ is also convex [19]. In the GCS literature, constraints are represented using sets, generally written $\mathcal{X}$, and are applied across the entire graph using the $\text{clh}(\mathcal{X})$ representation.
Figure 4: A set $\mathcal{S}\in\mathbb{R}^{2}$ (a) and the closure of the homogenization of $\mathcal{S}$ (b). The $\text{clh}(\mathcal{S})$ forms a cone in $\mathbb{R}^{3}$.
II-B Graph Theory
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{0},v_{1},\dots,v_{m}\}$ is the set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. When $\mathcal{E}$ consists of ordered pairs, the graph is directed. We will denote edges using ordered pairs $(v_{a},v_{b})$ to represent the directed connection from $v_{a}$ to $v_{b}$. Some equations apply to every directed edge in the graph. In this case, we will denote a generic edge in the graph as $e$,
A graph of convex sets is a directed graph in which each vertex is associated with a convex set. Formally, it is written as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{A}=\{A_{v_{0}},A_{v_{1}},...,A_{v_{m}}\}$ is a collection of convex sets, with each set $A_{v_{i}}$ assigned to vertex $v_{i}\in\mathcal{V}$. These convex sets may reside in arbitrary dimensions and can encode a wide variety of problem structures. Notably, the presence of an edge $(v_{a},v_{b})\in\mathcal{E}$ does not imply that the corresponding sets $A_{v_{a}}$ and $A_{v_{b}}$ intersect or overlap in $\mathbb{R}^{n}$.
Figure 5: A directed graph of convex sets.
Figure 5 shows an example graph of convex sets. The graph is directed, and each vertex is associated with a different convex shape in $\mathbb{R}^{2}$. In this illustration, the convex sets do not touch or overlap in $\mathbb{R}^{2}$.
II-C Graphs of Convex Sets (GCS) Optimization
The GCS optimization framework combines the strengths of continuous optimization and discrete graph searches. In this framework, a parameterized curve is assigned to each convex set. During optimization, curves corresponding to unvisited vertices are discarded, while curves at visited vertices connect at set boundaries to form a continuous trajectory. Unlike traditional graph searches, where the cost of an edge or vertex is a fixed value, the cost here varies continuously as the curves change within their respective convex sets. Minimizing the cost of traveling through the graph can be formulated using the objective function:
$\min_{x}\quad$ $\sum_{v\in\mathcal{V}}f_{v}(\vec{x}_{v})+\sum_{e\in\mathcal{E}}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (6a)
s.t.
$(\mathcal{V},\mathcal{E},\mathcal{A})\subseteq\mathcal{G}$ (6b)
$\vec{x}_{v}\in\mathcal{X}_{v}$ $\forall v\in\mathcal{V}$ (6c)
$\vec{x}_{a}\in\mathcal{X}_{e}$ $\forall e\in\mathcal{E}$ (6d)
$\vec{x}_{b}\in\mathcal{X}_{e}$ $\forall e\in\mathcal{E}.$ (6e)
Here, $x$ represents all the continuous variables defining the trajectory curve segments. The variables $\vec{x}_{v}$ are the design variables associated with the vertex $v$. The vertex cost function $f_{v}(\vec{x}_{v})$ represents the cost of the continuous variables within a vertex. The edge cost function $f_{e}(\vec{x}_{a},\vec{x}_{b})$ represents the cost of traversing an edge from $v_{a}$ to $v_{b}$, and couples the continuous variables of the connecting vertices. The vertices $\mathcal{V}$, edges $\mathcal{E}$, and associated convex sets $\mathcal{A}$ that form the optimal trajectory are a subset of the graph $\mathcal{G}$ that minimizes the overall cost. The sets $\mathcal{X}_{v}$ and $\mathcal{X}_{e}$ define the vertex and edge constraints that the design variables must satisfy.
A limitation of this formulation is that it does not provide a method to find $(\mathcal{V},\mathcal{E},\mathcal{A})\subseteq\mathcal{G}$ during optimization. To address this, the problem can be reformulated as a mixed integer program (MIP) by introducing binary variables to represent the vertices and edges used in the final solution. The binary variable $y_{v}\in\{0,1\}$ indicates whether a vertex is included in the solution, while the binary variable $y_{e}\in\{0,1\}$ indicates whether an edge is traversed.
The binary variables are multiplied into the cost and constraint terms of Equation 6. This ensures that the cost and constraints are active only when the vertex or edge is traversed, and the constraint is removed when the vertex or edge is not in the final solution. The MIP formulation is written as:
$\min_{x,y}\quad$ $\sum_{v\in\mathcal{V}}y_{v}f_{v}(\vec{x}_{v})$
$+$ $\sum_{e\in\mathcal{E}}y_{e}f_{e}(\vec{x}_{a},\vec{x}_{b})$ (7a)
s.t.
$y_{v}\in\{0,1\}$ (7b)
$y_{e}\in\{0,1\}$ (7c)
$y_{v}\vec{x}_{v}\in y_{v}\mathcal{X}_{v}$ $\forall v\in\mathcal{V}$ (7d)
$y_{e}\vec{x}_{a}\in y_{e}\mathcal{X}_{e}$ $\forall e\in\mathcal{E}$ (7e)
$y_{e}\vec{x}_{b}\in y_{e}\mathcal{X}_{e}$ $\forall e\in\mathcal{E}.$ (7f)
As the graph grows, the combinatorial nature of the binary variables makes this MIP intractable to solve directly. To address this, the binary variables are relaxed to continuous variables $y\in[0,1]$, and the problem is repeatedly solved using a branch-and-bound algorithm. In this formulation, the relaxed variables represent flow through the graph, from the start of the trajectory to its end. The flow variables are constrained to ensure flow conservation at all intermediate vertices. Although the problem remains NP-hard, this relaxation is tight and significantly improves tractability.
To handle the bilinear terms $y_{v}\vec{x}_{v}$ and $y_{e}\vec{x}_{a},y_{e}\vec{x}_{b}$, a change of variable is applied: $\vec{z}_{v}=y_{v}\vec{x}_{v}$ and $\vec{z}_{e,v}=y_{e}\vec{x}_{v}$. Assuming that $f_{v}$ and $f_{e}$ are convex functions and $\mathcal{X}_{v}$ and $\mathcal{X}_{e}$ are convex sets, this substitution transforms the problem into a mixed integer convex program (MICP), which drastically reduces the computation time per branch-and-bound iteration. The relaxed MICP can be written as:
$J(\vec{z},y)=$ $\sum_{v\in\mathcal{V}}y_{v}f_{v}\left(\frac{\vec{z}_{v}}{y_{v}}\right)+\sum_{e\in\mathcal{E}}y_{e}f_{e}\left(\frac{\vec{z}_{a}}{y_{a}},\frac{\vec{z}_{b}}{y_{b}}\right)$ (8a)
$\min_{z,y}$ $\quad J(\vec{z},y)$ (8b)
s.t.
$y_{v}\in[0,1]$ (8c)
$y_{e}\in[0,1]$ (8d)
$\sum y_{(a,v)}+\delta_{s}$
$=$ $\sum y_{(v,b)}+\delta_{f}=y_{v}\quad\forall v\in\mathcal{V}$ (8e)
$\vec{z}_{v}\in y_{v}\mathcal{X}_{v}\quad\forall v\in\mathcal{V}$ (8f)
$\vec{z}_{e,a}\in y_{e}\mathcal{X}_{e}\quad\forall e\in\mathcal{E}$ (8g)
$\vec{z}_{e,b}\in y_{e}\mathcal{X}_{e}\quad\forall e\in\mathcal{E}.$ (8h)
Equation (8e) introduces the flow conservation constraints. The variable $y_{(a,v)}$ denotes all the edge variables going to vertex $v$, and $y_{(v,b)}$ denotes all the edge variables leaving vertex $v$. The value $\delta_{s}=1$ if $v$ is the starting vertex and $\delta_{s}=0$ otherwise. The value $\delta_{f}=1$ if $v$ is the final vertex and $\delta_{f}$ is 0 otherwise. These create a source and sink within the graph at the vertices that contain the start and end of the trajectory. In other papers, the variables $z$ and $y$ are concatenated in Equations (8f),  (8g), and (8h), and the constraints are expressed using $\text{clh}(\mathcal{X})$. For clarity and consistency with the implementation in code, we omit this notation and leave the constraint as written here. For a geometrical interpretation of the resulting set, we refer the reader to Figure 4.
The previous GCS literature shows that the form of the MICP in Equation (8) can be used to find globally optimal trajectories. The convex relaxation is solved iteratively using a branch-and-bound algorithm, progressively adding constraints on the values of $y$ until the global optimal solution is found. We will leverage this formulation to find globally optimal collision-free trajectories.
II-D Generating The Graph Of Convex Sets
There are several ways to generate convex sets. In this work, we use the Iterative Regional Inflation by Semidefinite programming (IRIS) algorithm, as described in [29]. This algorithm is the standard approach for generating convex sets in previous GCS literature.
The IRIS algorithm works by first sampling a point in the configuration space. Then the algorithm iteratively adds halfspace planes to define a convex set, as well as increases the volume of an ellipsoid inscribed within the proposed convex set, until the volume is maximized. If a point is sampled within an obstacle, it is rejected, and the convex set is not created.
The IRIS algorithm can be configured to generate convex sets that either intersect or only touch at their boundaries. Once the free space is sufficiently sampled, each set is associated with a vertex in the graph, and the graph edges are created according to user-defined criteria. In this work, we generate sets that touch at their boundaries to avoid scenarios where the start and end points are contained within multiple sets. Sets that touch at their boundaries are connected to form the edges of the graph.
This method is advantageous because it can generate collision-free convex sets in dimensions higher than $\mathbb{R}^{3}$, such as the configuration space of a 7-DOF robot arm. However, its main drawback is that it requires sampling the configuration space and does not guarantee complete coverage of the free space. Consequently, trajectories computed using this graph are globally optimal with respect to the graph but not necessarily globally optimal with respect to the entire free space.
II-E Bézier Splines
Trajectories are often represented using piecewise polynomial functions known as splines. Many types of splines exist, such as B-splines, minimum volume splines, and Bézier splines. Previous GCS papers have used Bézier splines to represent an agent’s trajectory, and for consistency, we also use Bézier splines.
Bézier splines are computationally efficient to sample and modify. Their derivatives can be calculated easily as linear combinations of the control points defining the trajectory and are themselves lower-order splines. Because the trajectory passes through the first and last control points of a Bézier curve, continuity and differentiability constraints are intuitive to enforce in many applications.
The vector-valued equation for an $n^{\text{th}}$-order Bézier curve is
$\vec{B}(s)$ $=\sum_{k=0}^{n}\binom{n}{k}s^{k}(1-s)^{n-k}\vec{x}_{k},$ $0\leq s\leq 1,$ (9)
where $\vec{x}_{k}$ is the $k$th vector-valued control point of the curve and $\binom{n}{k}$ is the binomial coefficient. An $n^{\text{th}}$-order curve has $n+1$ control points. Individual Bézier curves are combined into a piecewise function to create a Bézier spline, which can be written as
$\vec{P}(r)$ $=\begin{cases}\vec{B}_{0}(r)&0\leq r<1\\ \vec{B}_{1}(r-1)&1\leq r<2\\ \vdots\\ \vec{B}_{n}(r-l)&l\leq r<l+1,\end{cases}$ (10)
where $r\in[0,n]$ is the parameter that interpolates along the spline and $l$ denotes the number of individual Bézier curve segments.
Bézier splines have a convex hull property [30], which ensures that a Bézier curve never leaves the convex hull of its control points (see Figure 6). In this illustration, the set of control points $C=\{x_{0},x_{1},x_{2},x_{3}\}$ defines a Bézier curve which never leaves $\text{conv}(C)$. This property is often used in collision avoidance algorithms to ensure safety.
In the GCS framework, a Bézier curve segment is assigned to each vertex of the graph. The control points are constrained to remain within or on the boundary of their associated closed convex set, ensuring that the spline can never leave the set. As the MICP problem is solved repeatedly, the curve segments eventually connect to form the final spline solution. Unused parts of the graph, where $y_{v}=y_{e}=0$, along with their associated Bézier curves, are discarded, leaving a single continuous spline traversing the graph of convex sets.
Figure 7a illustrates an unsolved GCS trajectory problem. The convex sets are depicted in various shades of grey, with the central obstacle shown in black. Stars mark the start and end positions, the individual Bézier curves are drawn in red, and small black circles indicate the Bézier control points. Figure 7b shows the solved trajectory, where the Bézier curve segments are connected to form a continuous path from the start to the goal, and unused curves are discarded.
Figure 6: The spline (shown in red) stays within the convex hull of its control points when $s\in[0,1]$.
Figure 7: Bézier curves associated with sets in GCS (a) and the solution to the GCS problem that connects Bézier curves to form a spline from start to end (b). Unused curves are removed.
III Methods
This section presents our methods for generating optimal collision-free trajectories using the GCS framework. We first describe how we generate convex sets to represent safe regions in the environment. We then outline a general strategy for formulating and implementing GCS-compatible constraints, addressing both vertex and edge constraints. We introduce the specific constraints used for trajectory planning in the space-time configuration space, including causality and kinematic constraints, as well as describe how obstacles are represented in the space-time configuration space. Finally, we conclude with the complete optimization formulation that enables time-dependent trajectories in 2D.
III-A GCS Constraints
The formulation and implementation of constraints, particularly those spanning edges in the graph, are among the most challenging aspects of the GCS framework. While prior work demonstrates that continuity and differentiability constraints can be implemented successfully within the GCS framework, the existing literature often lacks precise or accessible formulations of these constraints. Particularly, prior work shows results that must have implemented constraints that fulfill Equations (8g) and  (8h), but never show how this is done other than declaring the set $\mathcal{X}$ to represent the constraint. We aim to address this gap by providing detailed derivations of these constraints as well as providing a repeatable step-by-step strategy to apply constraints across graph edges.
Trivial constraints, such as continuity and differentiability of the spline, become non-trivial within the convex relaxation of the mixed integer program (see Equation (8)). The convex relaxation raises questions such as: what does it mean for two Bézier splines in adjacent vertices to be continuous when the edge between them is only partially active (e.g., when $y_{e}=0.5$)? What if multiple connecting edges are semi-active simultaneously? How can a spline segment connect to multiple segments at once?
Designing GCS-compatible edge constraints requires a different approach because constraints must adapt based on whether an edge is active, semi-active, or inactive. When an edge is active ($y_{e}=1$), the constraint must be fully enforced. For example, in spline continuity, this means the two curves in adjacent sets form a continuous piecewise function. When an edge is semi-active ($0<y_{e}<1$), the constraint must allow for a smooth transition between possible connections in the graph. Finally, when an edge is inactive ($y_{e}=0$), both sides of the constraint should become zero, effectively deactivating the constraint so it does not influence the design variables or their interactions across other edges. Designing constraints that satisfy these three behaviors while remaining continuous is a challenging task.
Our method for writing GCS-compatible edge constraints is summarized in Algorithm 1. The first step is to write the constraint as if it were active for a single edge in the solution subgraph. Then, both sides of the equation are multiplied by the decision variable that represents whether that edge is active. Next, the constraints of all edges flowing out of a single vertex are then summed together and simplified, resulting in a convex combination of potential constraints. Then, the bilinear products are replaced with the bilinear variable $z$. Finally, the constraint is repeated for each vertex in the graph, excluding the sink vertex.
Algorithm 1 Writing Edge Constraints for GCS
1:Write constraint $C$ for a single edge, in terms of the vertex control points $x_{v}$ and the connecting control points $\vec{x}_{b}$.
2:Multiply both sides of the constraint by the edge decision variable $y_{e}$.
3:Sum the constraints for all edges connected to the vertex $v$.
4:Simplify the expression using the relationship in Equation (8e) assuming that $\delta_{f}=0$, $y_{v}=\sum y_{(v,b)}$.
5:Substitute bilinear products with the variables $\vec{z}_{v}=y_{v}\vec{x}_{v}$ and $\vec{z}_{e,v}=y_{e}\vec{x}_{v}$.
6:Repeat the constraint for each vertex $v\in\mathcal{V},v\neq\text{sink}$.
The following subsections will give step-by-step derivations of the constraints that make up Equations 8f8g, and  8h.
III-B Convex Set Constraints
Like many other GCS applications, we rely on the convex hull property to guarantee safety at every point along the entire trajectory. Recall that each vertex in the graph is associated with a convex set, within which a Bézier curve segment is defined. Because a Bézier curve always remains within the convex hull of its control points, we can guarantee that the curve stays within a safe region if we constrain the control points to lie within the associated convex set. This constraint will be written in a way that is compatible with Equation 8f.
A polygonal convex set can be represented as the intersection of a finite collection of halfspaces. Let $\hat{n}_{i}$ denote the normal vector of the $i^{th}$ plane and let $p_{i}$ be a point lying somewhere on that plane. A generic point $\vec{r}$ lies on or behind the halfspace plane if $\hat{n}_{i}\cdot\vec{r}\leq\hat{n}_{i}\cdot p_{i}$. By stacking each face normal as a row vector in a matrix $A$, and stacking the corresponding dot products $\hat{n}_{i}\cdot p_{i}$ into a vector $\vec{d}$, we can determine whether a point $\vec{r}$ lies within the convex set by checking the linear inequalities formed by $A\vec{r}\leq\vec{d}$.
The variable $\vec{x}_{v,i}$ represents the $i^{th}$ control point in the curve associated with the vertex $v$. The linear inequality $A_{v}\vec{x}_{v,i}\leq\vec{d}_{v}$ is applied to each control point in every set. Rewriting this in terms of the bilinear variable $\vec{z}_{v,i}=y_{v}\vec{x}_{v,i}$ we get the equation
$A_{v}\frac{\vec{z}_{v,i}}{y_{v}}\leq\vec{d}_{v}$ (11)
which can be rearranged to
$A_{v}\vec{z}_{v,i}\leq\vec{d}_{v}y_{v}.$ (12)
III-C Continuity Constraints
We begin by deriving the continuity constraints using Algorithm 1. Our goal is for the last control point of a curve to match the first control point of the connecting curve when the edge is fully active. Additionally, the constraint should allow for continuous transitions between possible connections when edges are only partially active. Because this constraint requires variables that are related by the edges in the graph, the following constraint formulation will fall under the category of Equations  8g and  8h.
We start by writing the constraint for a single generic edge in the graph:
$\vec{x}_{v,n}=\vec{x}_{b,0}.$ (13)
Here, $\vec{x}_{v,n}$ denotes the last control point of the current curve in set $v$, and $\vec{x}_{b,0}$ denotes the first control point of the connecting set $b$.
We then multiply both sides of the equation by the edge decision variable connecting set $v$ with set $b$, $y_{(v,b)}$, making the constraint fully active when the edge is included in the solution and inactive otherwise:
$y_{(v,b)}\vec{x}_{v,n}=y_{(v,b)}\vec{x}_{b,0}.$ (14)
Next, we sum these constraints over all edges connected to vertex $v$:
$\sum_{b}y_{(v,b)}\vec{x}_{v,n}=\sum_{b}y_{(v,b)}\vec{x}_{b,0}.$ (15)
Assuming the flow conservation constraint from Equation 8e holds, the left-hand side simplifies to:
$y_{v}\vec{x}_{v,n}=\sum_{b}y_{(v,b)}\vec{x}_{b,0}.$ (16)
Figure 8: A vertex $v$ with adjacent vertices $b_{i}$. The continuity constraint ensures that the final Bézier control point falls in the convex hull of the connecting edges’ first Bézier control point (a), and then connects two edges as $y_{v,b_{i}}\rightarrow 1$ (b).
Referring back to Equations (3) and (4), this constraint ensures that the last control point of a set lies within the convex hull of the first control points of the connected sets, weighted by their corresponding decision variables $y_{(v,b)}$. When applied with the constraint requiring all control points to remain within their respective convex sets, it ensures that the control points meet at the boundary between sets when $y=1$, as that is the only place where both constraints can be met simultaneously. This enforces continuity between curve segments when edges are active and enables smooth transitions in the convex relaxation. If all edges and vertices are inactive, the constraint has no effect. Because this formulation relates a vertex to its outgoing edges, it is applied to every vertex except the sink vertex, which is implicitly captured through the other edge constraints in the graph. Alternatively, the constraint could be formulated using ingoing edges, in which case the source vertex would be excluded instead. Figure 8 illustrates this constraint.
We then substitute the bilinear variables, $\vec{z}_{v,n}=y_{v}\vec{x}_{v,n}$ and $\vec{z}_{(v,b),0}=y_{(v,b)}\vec{x}_{b,0}$, to rewrite the constraint as:
$\vec{z}_{v,n}$ $=\sum_{i=0}^{n}\vec{z}_{(v,b),0}.$ (17)
III-D Differentiability Constraints
Next, we derive the differentiability constraint to ensure that the spline is differentiable with respect to the interpolation variable $r$ at the connection point between curves (see Equation (10)). Similar to the continuity constraint, the differentiability constraint will fall under the category of Equations  8g and  8h.
For the path to be differentiable where the two curves connect, the vector between the last two control points of one curve must equal the vector between the first two control points of the connecting curve. For the edge $(v,b)$, this constraint can be written as:
$\vec{x}_{v,n}-\vec{x}_{v,n-1}=\vec{x}_{b,1}-\vec{x}_{b,0}.$ (18)
As before, we multiply both sides of the equation by the edge decision variable $y_{(v,b)}$, making the constraint active when the edge is in the solution. The equation can be written as:
$y_{(v,b)}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (19)
We then sum these constraints over all edges connected to vertex $v$:
$\sum_{b}y_{(v,b)}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=\sum_{b}y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (20)
Assuming the flow conservation constraint from Equation 8e holds, the left-hand side simplifies to:
$y_{v}(\vec{x}_{v,n}-\vec{x}_{v,n-1})=\sum_{b}y_{(v,b)}(\vec{x}_{b,1}-\vec{x}_{b,0}).$ (21)
We then substitute the bilinear variables, $\vec{z}_{v,0}=y_{v}\vec{x}_{v,n}$, $\vec{z}_{v,n-1}=y_{v}\vec{x}_{v,n-1}$, $\vec{z}_{(v,b),1}=y_{(v,b)}\vec{x}_{b,1}$, and $\vec{z}_{(v,b),0}=y_{(v,b)}\vec{x}_{b,0}$, to rewrite the constraint as:
$\vec{z}_{v,n}-\vec{z}_{v,n-1}$ $=\sum_{b}(\vec{z}_{(v,b),1}-\vec{z}_{(v,b),0}).$ (22)
However, this formulation is inefficient because it requires defining two $z$ variables for every directed edge, resulting in $|\mathcal{E}|\times 2$ extra variables to enforce this constraint, where $|\mathcal{E}|$ represents the cardinality of the edges of the graph. To reduce the number of variables, we introduce a final substitution variable, $\vec{z}_{(v,b_{i}),\text{diff}}=\vec{z}_{(v,b_{i}),1}-\vec{z}_{(v,b_{i})}$. This allows us to express the differentiability constraint as:
$(\vec{z}_{v,n}-\vec{z}_{v,n-1})=\sum_{i=0}^{n}\vec{z}_{(v,b_{i}),\text{diff}}.$ (23)
This formulation only requires $|\mathcal{E}|$ additional variables to enforce differentiability, which accelerates the optimization process.
III-E Space-Time Configuration Constraints
We now introduce the space-time configuration constraints, which are essential for generating realistic trajectories in the space-time configuration space. In the space-time configuration space, it is possible to construct splines that are non-causal and result in infinite velocities. To prevent this, we impose a constraint that will enforce causality and limit the maximum velocity on the spline. This constraint will be written in a way that is compatible with Equation 8f
The maximum velocity of the spline $v_{\text{spline}}$ between control points can be bounded by the changes in the $x$, $y$, and $t$ components of the control points, so that
$v_{\text{spline}}\leq\frac{\lVert(\vec{x}_{v,i+1}-\vec{x}_{v,i})\rVert_{xy,2}}{(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}}$ $\leq v_{\text{max}}$ (24)
$(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ $\geq 0$ (25)
where $v_{\text{max}}$ is the maximum allowable velocity. The notation $\lVert\cdot\rVert_{xy,2}$ denotes the 2-norm with respect to just the $x$ and $y$ components of the control points, and $(\cdot)_{t}$ is the value of the time component within the parentheses.
Figure 9: The velocity constraint forms a cone in $\mathbb{R}^{3}$. Each control point must lie within the cone formed by the previous control point to ensure the slope remains below the maximum velocity.
Figure 10: The intersection of all velocity cone constraints remains convex (a). The final control point must lie within this intersection, which is equivalent to just placing a velocity cone around the second-to-last control point (b).
Figure 11: A dynamic obstacle is shown at different times in $\mathbb{R}^{2}$ (a). The initial and final positions of the vertices are put in the space-time configuration space (b). The vertices are connected to form a shape showing its trajectory in $\mathbb{R}^{3}$ (c).
To make this constraint compatible with a convex solver, we can rearrange the constraint as:
$\lVert(\vec{x}_{v,i+1}-\vec{x}_{v,i})\rVert_{xy,2}$ $\leq v_{\text{max}}(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ (26)
$(\vec{x}_{v,i+1}-\vec{x}_{v,i})_{t}$ $\geq 0$ (27)
Substituting the bilinear variable $\vec{z}_{v,i}$, this becomes:
$\left\lVert\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)\right\rVert_{xy,2}$ $\leq v_{\text{max}}\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)_{t}$ (28)
$\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)_{t}$ $\geq 0.$ (29)
The following constraint
$\left(\frac{\vec{z}_{v,i+1}}{y_{v}}-\frac{\vec{z}_{v,i}}{y_{v}}\right)_{t}\geq\epsilon$ (30)
can be substituted to ensure that control points maintain a minimum time separation.
Finally, simplifying $y_{v}$ yields the constraint in its final form:
$\lVert(\vec{z}_{v,i+1}-\vec{z}_{v,i})\rVert_{xy,2}$ $\leq v_{\text{max}}(\vec{z}_{v,i+1}-\vec{z}_{v,i})_{t}$ (31)
$(\vec{z}_{v,i+1}-\vec{z}_{v,i})_{t}$ $\geq y_{v}\epsilon.$ (32)
The geometric intuition of this constraint is that it forms a cone in $\mathbb{R}^{3}$, depicted in Figure 9a. Due to the convex hull property, the instantaneous slope of the Bézier curve never exceeds the slope made between two adjacent control points. Therefore, if the slope between two control points is less than the maximum velocity, the curve will never violate the constraint. To ensure that the maximum velocity constraint is satisfied, we must ensure that connecting adjacent control points is within the cone formed by the previous control point. This is shown in Figure 9.
While it is clear that this constraint is convex for any individual pair of adjacent control points, it may not be immediately obvious whether the combined constraints for all adjacent pairs remain convex. Figure 10 illustrates that the feasible region for the final control point is the intersection of all the velocity cone constraints. This intersection of feasible regions is equivalent to just taking the velocity constraint at the previous control point.
III-F Space-Time Obstacles
Moving obstacles in $\mathbb{R}^{2}$ can be represented in the space-time configuration space by concatenating the initial positions of their vertices with the initial time, then concatenating the final positions with the final time, and finally connecting all vertices to form the obstacle in the space-time $\mathbb{R}^{3}$ configuration. Figure 11 illustrates this process. This method assumes that obstacles are either static or move at constant velocity. More complex motions require finer sampling and will only approximate their true motion in $\mathbb{R}^{3}$.
III-G Cost
The final component needed to implement the space-time GCS formulation is the cost function. It is important to choose the vertex cost function $f_{v}$ and the edge cost function $f_{e}$ so that the overall objective remains convex and simplifies when decision and relaxation variables are introduced, as shown in Equation 8. In this work, we aim to minimize the distance traveled by the trajectory in the $(x,y)$ plane. A well-known property of Bèzier curves is that the sum of the distances between adjacent control points is an upper bound on the curve’s total length [31]. For simplicity, we assume that all convex sets touch in $\mathbb{R}^{n}$ and that there is no additional cost for traversing an edge, so that $f_{e}(x_{e})=0$. Thus, the cost function for each node can be written in terms of the design variables as:
$f_{v}(x_{v})=\sum_{i=0}^{n-1}\lVert x_{v,i+1}-x_{v,i}\rVert_{xy,2}.$ (33)
Substituting $z_{v,i}=y_{v}x_{v,i}$ into $f_{v}$ we get
$f_{v}\left(\frac{z_{v}}{y_{v}}\right)=$ $\sum_{i=0}^{n-1}\left\lVert\frac{z_{v,i+1}}{y_{v}}-\frac{z_{v,i}}{{y_{v}}}\right\rVert_{xy,2}$ (34)
$=$ $\frac{1}{\lvert y_{v}\rvert}\sum_{i=0}^{n-1}\rVert z_{v,i+1}-z_{v,i}\rVert_{xy,2}.$
We can remove $y_{v}$ from the norm and, because it is non-negative, drop the absolute sign around it. The total cost function can be written as:
$J(y,z)=$ $\sum_{v\in\mathcal{V}}y_{v}\frac{1}{y_{v}}\sum_{i=0}^{n-1}\lVert z_{v,i+1}-z_{v,i}\rVert_{xy,2}$ (35)
$J(z)=$ $\sum_{v\in\mathcal{V}}\sum_{i=0}^{n-1}\rVert z_{v,i+1}-z_{v,i}\rVert_{xy,2}.$
This is the cost function that we used to generate collision-free trajectories using the GCS formulation.
III-H Optimization Formulation
We incorporate the cost function along with the edge and vertex constraints into the overall optimization framework described by Equation 8. The resulting formulation is a relaxed mixed-integer convex program (MICP), written as:
$J(z)=$ $\sum_{v\in\mathcal{V}}\sum_{i=0}^{n-1}\rVert z_{v,i+1}-z_{v,i}\rVert_{xy,2}$ (36a)
$\min_{z,y}$ $\quad J(z)$ (36b)
s.t.
$y_{v}\in\{0,1\}$ (36c)
$y_{e}\in\{0,1\}$ (36d)
$\sum y_{(a,v)}+\delta_{s}$
$=\sum y_{(v,b)}+\delta_{f}=y_{v}\quad\forall v\in\mathcal{V}$ (36e)
$A_{v}z_{v,i}\leq b_{v}y_{v}.$ (36f)
$z_{v,n}=\sum_{b}\vec{z}_{(v,b),0}\quad\forall v\in\mathcal{V}$ (36g)
$z_{v,n}{-}z_{v,n-1}=\sum_{b}z_{(v,b),\text{diff}}\quad\forall v\in\mathcal{V}$ (36h)
$\lVert(z_{v,i+1}{-}z_{v,i})\rVert_{xy,2}\leq$
$v_{\text{max}}(z_{v,i+1}-z_{v,i})_{t}$ (36i)
$(\vec{z}_{v,i+1}-\vec{z}_{v,i})_{t}\geq y_{v}\epsilon.$ (36j)
This final formulation includes relaxed binary variables $y$ for each vertex and edge, $n$ control point vectors $z_{v,i}$ within each vertex, and additional vectors representing continuity $\vec{z}_{(v,b),0}$ and differentiability $z_{(v,b),\text{diff}}$ for each directed edge. Each of these functions and constraints is convex and is compatible with general convex solvers.
IV Results
In this section, we apply our space-time GCS formulation to a range of scenarios and evaluate performance. First, we will validate our formulation against the vanilla GCS formulation in a scenario with only static obstacles. Then, we demonstrate its effectiveness in cluttered environments with containing many dynamic obstacles.
The simulations and algorithms were implemented in Python and executed on an AMD Ryzen 7 9700X processor with 32 GB of RAM. The convex optimization problems were formulated using CVXPY and solved with the open-source solver CLARABELL.
IV-A Standard GCS Comparison
Because the standard GCS formulation is not compatible with dynamic environments, we will compare the space-time method against it in a static scenario. We demonstrate that it produces the same optimal trajectory in this setting. The environment is decomposed into a graph of convex sets in two dimensions for the standard GCS formulation, and three dimensions for the space-time formulation.
The scenario parameters are as follows. An agent with state (x,y,t) begins at ($0.5m$, $0m$, $0s$) and must move to ($0.5m$, $1m$, $1s$). The agent’s maximum velocity is $2m/s$. A static obstacle is present with vertices at ($0.3m$, $0.2m$), ($0.6m$, $0.2m$), ($0.6m$, $0.4m$), and ($0.3m$, $0.4m$). Because this obstacle is slightly off center, there is a single geometric minimum distance that avoids the obstacle between the start and end points, which is $1.0318m$. Figure 12 depicts the scenario in 2d and 3d. The static obstacle is shown in black, with its vertices aligned vertically between $t=0s$ and $t=1s$. The start and end points are shown in green and red, respectively.
Figure 12: The environment shown in 2D (a) and an isometric view of the environment that will be used in the space-time GCS formulation (b).
Figure 13 shows the trajectory generated by the standard GCS formulation. The sets and their associated Bèzier curve control points are color-coded accordingly. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.2s$. The optimal distance traveled is $1.0318m$, matching the geometric minimum.
Figure 13: Path generated by the standard GCS formulation.
Figure 14 shows the control points generated by the space-time method from a top-down and an isometric view. When viewed from above, the geometric path is identical to the standard GCS solution. The isometric view illustrates the vertical placement of each spline control point over time. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.29s$. The optimal distance traveled in 2d is also $1.0318m$, which matches both the standard GCS solution and the true geometric minimum. These results demonstrate that our modified GCS formulation produces trajectories identical to vanilla GCS in static environments, validating the correctness of our implementation.
Figure 14: Resulting sets and trajectory of the space-time GCS shown from the top (a) and in an isometric view (b).
IV-B Dynamic Environment
Figure 15: Resulting sets and collision-free trajectory of the space-time GCS shown from the top (a) and in an isometric view (b).
Next, we present a simple scenario with a single moving obstacle. In this case, multiple optimal trajectories have the same global minimum distance. We show that our method finds an optimal trajectory through the graph of convex sets while avoiding the moving obstacle. It should be noted that the standard GCS method can’t be used to solve this scenario without introducing a non-linear solver to determine the timing of the path and avoid collisions.
The scenario parameters are as follows. An agent with state $(x,y,t)$ starts at ($0.5m$, $0m$, $0s$) and must move to ($0.5m$, $1m$, $1s$). The agent’s maximum velocity is $2m/s$. There is a square-shaped dynamic obstacle with sides of length $0.2m$. The square starts centered at ($0m$, $0.5m$, $0s$) and moves to ($1m$, $0.5m$, $1s$). The geometric minimum distance avoiding the obstacle from the agent’s start location and end location is $1.0m$.
Figure 15 shows the control points generated by our method from a top-down and isometric view. When viewed from above, the geometric path moves in a straight line from start to finish. The isometric view shows that the agent moves quickly in the beginning, as evidenced by the trajectory’s slope, to move past the obstacle, then slows down. The graph of convex sets consists of four sets and eight directed edges. The total computation time was $0.52s$. The optimal distance traveled in 2D is also $1.0m$, which is the true geometric minimum distance. This example demonstrates the capability of our approach in generating collision-free, optimal trajectories in the presence of a dynamic obstacle, without relying on an initial guess.
IV-C Cluttered Environment
The previous example shows that our method can find optimal trajectories avoiding a single obstacle. However, in practical scenarios, there are often significantly more obstacles to navigate. This next scenario illustrates how our method performs in a cluttered environment, finding optimal paths through the graphs of convex sets, and presents a heuristic for convex set generation that improves results. We also discuss some of the shortcomings that arise due to the method used to generate the convex sets.
The scenario parameters are as follows. An agent starts at ($0.5m$, $0m$, $0s$) and must move to ($0.5m$, $1m$, $1s$). The agent’s maximum velocity is $3m/s$, allowing it to travel up to three times the distance from start to finish in the allotted $1s$. Twenty obstacles are randomly initialized between ($0m$, $0.2m$) and ($1m$, $0.8m$) with random constant velocities. Figure 16 shows the starting and ending 2D cross sections of this dynamic environment. Figure 17 shows what this looks like in 3D from the side and an isometric view. Finding an initial guess for a trajectory that avoids all the obstacles in this environment would be extremely challenging, if not impossible.
Figure 16: The obstacles move at constant velocity from a random start location to a random end location. Arrows indicate the direction of motion. The start and end of the trajectory are shown in green and red, respectively.
Figure 17: Cluttered environment shown from the top (a) and side view (b).
The free space is decomposed into a graph of convex sets in 3D. The IRIS splitting algorithm generates each convex set by expanding from a seed point to create the largest possible collision-free convex set containing the seed point. The algorithm is seeded by randomly sampling the environment, discarding samples that fall within obstacles or existing sets, and then expanding the graph of convex sets. Figure 18a shows a 3D view of the graph of convex sets after sampling 80 random points and appropriately discarding the samples. In 3D, many of the sets are visually obscured by obstacles or other sets. Figure 18b shows a cross-section of the graph of convex sets at ($t=0.80s$), with sets color-coded and obstacles shown in black. With this random seed, the graph of convex sets consists of fourteen sets and nine directed edges.
Figure 18: 3D view of the graph of convex sets after sampling 80 points (a) and a cross section of the graph at $t=0.80s$ (b).
Figure 19 shows a trajectory generated by our method. There are no collisions with any of the 20 obstacles in the environment. However, while the trajectory is globally optimal with respect to the graph of convex sets, the sets themselves do not provide complete coverage of the free space. As seen in Figure 18b, there is a low density of sets in areas with many obstacles. This results in a trajectory that is not globally optimal with respect to the geometric minimum distance. Figure 19 shows the trajectory in 3D from the top and an isometric view. The total computation time was $4.12s$. The optimal distance traveled is $1.28m$.
Figure 19: A top view of the trajectory generated by our method (a) and a perspective view of the trajectory (b).
We ran the same scenario as above while varying the number of IRIS samples, repeating the experiment 100 times with different seeds. For each run, we recorded the number of sets and edges created in the graph, the final trajectory cost, and the total computation time required to generate the graph and optimize the trajectory. As expected, when the number of samples increased, both the number of sets and edges grew, resulting in a lower average trajectory cost.
Table I summarizes how the average results change with the number of samples used to seed the IRIS algorithm. Computation time increases with the size of the graph, while the distance of the trajectory decreases. The time required to compute the trajectory grows substantially as the graph expands. While some GCS papers propose strategies such as pruning the graph after the first branch-and-bound iteration to reduce computation time, we did not implement these methods in this work. Figure 20 shows an example cross section at $t=0.8s$ after sampling 1000 points to seed the IRIS algorithm, resulting in denser set coverage through the center of the environment and enabling solutions closer to the geometric minimum.
TABLE I: Mean Effect of IRIS Sampling on GCS Trajectory Optimization
Samples Sets Edges Cost (m) Time (s)
80 15.22 22.71 1.24 4.75
100 17.05 26.45 1.23 5.28
250 27.58 52.28 1.12 11.50
500 40.16 95.27 1.05 26.80
1000 55.62 148.44 1.03 82.30
Figure 20: Cross section $t=0.8s$ with 1000 samples used to seed the IRIS algorithm (a). The convex sets are densely packed in the center of the environment, resulting in a shorter trajectory (b).
While the method successfully avoids all obstacles, these results underscore the impact of convex set density on trajectory optimality. Sparse coverage in regions with many obstacles can lead to suboptimal paths, even if the solution is globally optimal within the constructed graph. Increasing the number of samples improves set density and solution quality but comes at the cost of significantly increased computation time.
V Conclusion
We have presented a GCS formulation for trajectory generation in cluttered, dynamic environments. The constraint development strategy introduced here extends the applicability of GCS to a broader range of trajectory optimization problems. Although the convex set generation method limits solution quality, the GCS framework provides safety and optimality guarantees that are rarely available in other trajectory optimization methods. Additionally, the space-time formulation enables GCS to handle dynamic environments effectively. Notably, this approach allows optimal trajectories to be generated without requiring a good initial spatial or temporal guess. Overall, this work demonstrates a strategy to implement GCS constraints, as well as how the GCS framework can be extended to handle dynamic environments.
Acknowledgment
This project was supported by NSF SBIR Phase 2 Award Number 2404858 and 4D Avionic Systems, LLC. We thank Dr. Garth Thompson for his feedback and insightful suggestions.